Neelon Corporation has two divisions: Southern Division and Northern Division. The following data are for the most recent operating period:

Total Company Southern Division Northern Division
Sales $ 418,000 $ 193,000 $ 225,000
Variable expenses $ 130,880 $ 79,130 $ 51,750
Traceable fixed expenses $ 186,000 $ 77,000 $ 109,000
Common fixed expense $ 79,420 $ 36,670 $ 42,750

The common fixed expenses have been allocated to the divisions on the basis of sales.

What is the company's overall net operating income if it operates at the break-even points for its two divisions?

  • $21,700

  • $(79,420)

  • $0

  • $(265,420)
Answer #1

Contribution margin of southern division = Sales - Variable expenses

= 193,000 - 79,130

= $113,870

Contribution margin ratio of southern division = Contribution margin/sales

= 113,870/193,000 = 59%

Contribution margin of northern division = Sales - Variable expenses

= 225,000 - 51,750

= $173,250

Contribution margin ratio of northern division = Contribution margin/sales

= 173,000/225,000 = 77%

Break even sales of southern division = Fixed costs/Contribution margin ratio

= 77,000/59%

= $130,508

Break even sales of northern division = Fixed costs/Contribution margin ratio

= 109,000/77%

= $141,558

Income statement (At break even points of two division)

Sales

- Variable expenses

Southern division

130,508

- 53,508

Northern division

141,558

- 32,558

Total company

272,066

- 86,066

Contribution margin   

-Traceable fixed cost

77,000    

-77,000

109,000      

- 109,000   

186,000

- 186,000

Segment margin 0    0    0
-Common fixed costs - 79,420
Net income - 79,420

Correct option is (b) i.e. (79,420)

Please do comment if you have any query, thanks.